1. A strip of magnesium of mass 15g is dropped into an open beaker of dilute hydrochloric acid. Calculate the work done by the system as a result of reaction. The atmosphere pressure is 1atm and temperature is 25°C. Also, calculate the work done if the reaction is carried out in closed beaker.


1.       A strip of magnesium of mass 15g is dropped into an open beaker of dilute hydrochloric acid. Calculate the work done by the system as a result of reaction. The atmosphere pressure is 1atm and temperature is 25°C. Also, calculate the work done if the reaction is carried out in closed beaker.  

Grade:11

1 Answers

Leah hanford
6 years ago
Mg + 2HCl-----> MgCl2 + H2
n=(15 g magnesium)*(1mol/24.305g)*(1mol H2/1mol Mg)=0.6171569636 mol H2
R=0.082054 atmL/molK
T=25+273.15=298.15K
P= 1 atm
PV=nRT
V=nRT/P=(0.6171569636 mol )(0.082054 atmL/molK)(298.15K)/1 atm=15.09837488 L
W=-PV=- (1 atm)(15.09837488 L)= -15.09837488 atmL= 1.529918327 kJ